# Omar I. Al-Bataineh

### 27 Dec 2023 v1

@@ Abstract
This paper describes a formal general-purpose automated program repair (APR) framework based on the concept of program invariants. In the presented repair framework, the execution traces of a defected program are dynamically analyzed to infer specifications $\varphi_{correct}$ and $\varphi_{violated}$, where $\varphi_{correct}$ represents the set of likely invariants (good patterns) required for a run to be successful and $\varphi_{violated}$ represents the set of likely suspicious invariants (bad patterns) that result in the bug in the defected program. These specifications are then refined using rigorous program analysis techniques, which are also used to drive the repair process towards feasible patches and assess the correctness of generated patches. We demonstrate the usefulness of leveraging invariants in APR by developing an invariant-based repair system for performance bugs. The initial analysis shows the effectiveness of invariant-based APR in handling performance bugs by producing patches that ensure program’s efficiency increase without adversely impacting its functionality.
@@ 2 Invariant-based Program Repair Framework
It is important to categorize and distinguish inferred patterns (invariants) into good and bad patterns, especially when dealing with programs that have several functional requirements. This helps to identify the set of desired invariants to be maintained and violated invariants to be repaired when modifying the code. It also helps to identify the set of relevant invariants for the particular bug being analyzed. The soundness of inferred $\varphi_{correct}$ and $\varphi_{violated}$ depends heavily on the soundness of the employed invariant inference tool as well as the invariant refinement process. Increasing the amount of program behavior exercised using reachability analysis increases the likelihood that $\varphi_{correct}$ and $\varphi_{violated}$ are true.
@@ 4 Related Work
Patch overfitting in APR: Several solutions have been developed to alleviate the overfitting problem in APR, such as symbolic specification inference [8], machine learning-based prioritization of patches [1], fuzzing-based test-suite augmentation [5], and concolic path exploration [12]. These solutions rely on limited incomplete test cases and do not guarantee the general correctness of the patches. Compared to those approaches that generate test inputs, invariant-based APR automatically generates and refines desired invariants that need to be maintained and violated invariants that need to be repaired when modifying the code, which makes the approach more reliable than existing repair approaches.

### 26 Jan 2024 v2

@@ Abstract
This paper describes a formal general-purpose automated program repair (APR) framework based on the concept of program invariants. In the presented repair framework, the execution traces of a defected program are dynamically analyzed to infer specifications $\varphi_{correct}$ and $\varphi_{violated}$, where $\varphi_{correct}$ represents the set of likely invariants (good patterns) required for a run to be successful and $\varphi_{violated}$ represents the set of likely suspicious invariants (bad patterns) that result in the bug in the defected program. These specifications are then refined using rigorous program analysis techniques, which are also used to drive the repair process towards feasible patches and assess the correctness of generated patches. We demonstrate the usefulness of leveraging invariants in APR by developing an invariant-based repair system for performance bugs. The initial analysis shows the effectiveness of invariant-based APR in handling performance bugs by producing patches that ensure program’s efficiency improvement without adversely impacting its functionality.
@@ 2 Invariant-based Program Repair Framework
It is important to categorize and distinguish inferred patterns (invariants) into good and bad patterns, especially when dealing with programs that have several functional requirements. This helps to identify the set of desired invariants to be maintained and violated invariants to be repaired when modifying code. It also helps to identify the set of invariants that are relevant to the analyzed bug. The soundness of inferred $\varphi_{correct}$ and $\varphi_{violated}$ depends heavily on the soundness of the employed invariant inference tool as well as the invariant refinement process. Increasing the amount of program behavior exercised using reachability analysis increases the likelihood that $\varphi_{correct}$ and $\varphi_{violated}$ are true.
@@ 4 Related Work
Patch overfitting in APR: Several solutions have been developed to alleviate the overfitting problem in APR, such as symbolic specification inference [8], machine learning-based prioritization of patches [1], fuzzing-based test-suite augmentation [5], and concolic path exploration [12]. These solutions rely on limited incomplete test cases and do not guarantee the general correctness of the patches. Compared to those approaches that generate test inputs, invariant-based APR automatically generates and refines desired invariants that need to be maintained and violated invariants that need to be repaired when modifying code, which makes the approach more reliable than existing repair approaches.
@@ 5 Conclusion and Future Work
Acknowledgements
The author would like to thank Leon Moonen for his insightful discussion on the idea of the paper. This work is supported by the Research Council of Norway through the secureIT project (IKTPLUSS #288787).